Find the equation of the line that passes through (5, -3) and is parallel to the line 3x + 4y = 8.
Answer by Gogonati(809)

Since the lines are parallel their slopes are equal: y=-3x/4+2 m=-3/4
The equation of the line that pases at point (5,-3) with slope -3/4 is:
y+3=-3/4(x-5) or 3x+4y=3.